OpenStudy (anonymous):

How do the graphs of y =1/x and y = 5/x +6 compare?

jimthompson5910 (jim_thompson5910):

If you take 1/x and stretch it vertically by a factor of 5, you'll get y = 5/x If you shift y = 5/x six units to the left, you'll get y = 5/(x+6)

jimthompson5910 (jim_thompson5910):

so you can see how they're related since they are essentially the same shape...one is just has been stretched and shifted